OpenStudy (anonymous):

jack usually mows his lawn in 4 hours. marilyn can mow the same yard in 7 hours. How much time would it take for them to mow the lawn together. Can someone please help me with this..think I am overthinking and can not get it!!!

OpenStudy (saifoo.khan):

28/11 hrs.

OpenStudy (saifoo.khan):

Nope, \[\frac{28}{11}\]

OpenStudy (saifoo.khan):

NO problem. Have more? =D

OpenStudy (saifoo.khan):

Sure! =D They are easy!

OpenStudy (saifoo.khan):

Yes, Always! xD

OpenStudy (saifoo.khan):

Man!!! i hate this question! literally! lol..

OpenStudy (saifoo.khan):

just a sec, let me solve..

OpenStudy (anonymous):

i hate these questions to.....not any good at word problems

OpenStudy (saifoo.khan):

k = 1/18 d = k*r^2 d = 1/18 * (70)^2 d = 2450/9 d = 272.22

OpenStudy (anonymous):

so would this be 272.2 feet?

OpenStudy (saifoo.khan):

It can be marked as wrong, what does mathlab says about the input format?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

simplify your answer as an integer or a fraction

OpenStudy (saifoo.khan):

Then go for 2450/9

OpenStudy (anonymous):

oh man...you are awesome!!!!!!! Wish I had a direct line to you...LOL

OpenStudy (saifoo.khan):

LOl, are in UoP?

OpenStudy (anonymous):

yes...and I have tried everything......with those two questions and just could not comprehend them for some reason

OpenStudy (saifoo.khan):

Cool. no worries.. =D

OpenStudy (saifoo.khan):

lol, that happens. actually im a private tutor for UoP students who are stuck with their math class..

OpenStudy (saifoo.khan):

No doubt! =D

OpenStudy (saifoo.khan):

Are u on fb?

OpenStudy (saifoo.khan):

http://www.facebook.com/saifoo.khan

OpenStudy (anonymous):

ok cool...thank you so much for your help

OpenStudy (saifoo.khan):

np!!

OpenStudy (saifoo.khan):

umm, can we discuss this on fb? if u dont mind?

OpenStudy (saifoo.khan):

=)